# Temperature Dependent Mechanical and Structural Properties of Uniaxially Strained Planar Graphene

### v1

@@ Abstract
Using molecular dynamics simulations in a planar graphene sheet, we investigate the temperature dependence of its mechanical behavior under uniaxial tensile stress applied either along the armchair or the zigzag direction. Stress-strain curves are calculated for different temperatures and the corresponding dependence of various elastic parameters, like the Young modulus, the third-order elastic modulus, the tensile strength and failure strain, is presented. Fracture stress and strain, as well as the Young modulus, decrease almost linearly with temperature. The distributions of bond lengths and bond angles at different strains and temperatures are also discussed and approximate analytical expressions are presented. The latter describe accurately the numerically obtained distributions.
@@ 1 Introduction
Since the discovery of graphene, there have been a number of investigations into its mechanical behavior. Despite the difficulty of applying controlled mechanical loads at the nanoscale, experimental studies have verified an exceptional value of stiffness and extremely high tensile strength ref-Lee-Science-2008 ; ref-Cao-NatComm-2020 ; ref-Varillas-MechSci-2024 , in accordance with corresponding theoretical predictions. There exist a number of related numerical investigations using molecular dynamics (MD) simulations with a variety of potential functions ref-Zakharchenko-PRL-2009 ; ref-Zhao-NanoLet-2009 ; ref-Neek-Amal-PRB-2010 ; ref-Shen-MatDes-2010 ; ref-Tsai-MatDes-2010 ; ref-Zhao-JAP-2010 ; ref-Wei-JCP-2011 ; ref-Dmitriev-CompMatSci-2012 ; ref-Kalosakas-JAP-2013 ; ref-Sgouros-2DMat-2016 ; ref-Souros-SciRep-2018 ; ref-Lebedeva-PhysE-2019 ; ref-Savin-MechofMat-2019 ; ref-Aparicio-Nanotech-2021 ; ref-deSousa-Intech-2021 ; ref-Kalosakas-Mat-2021 ; ref-Flores-PhysChem-2022 ; ref-Fthenakis-FrontChem-2022 ; ref-Pacheco-Sanjuan-Carbon-2023 ; ref-Sgouros-PRE-2024 , density functional theory ref-Kudin-PRB-2001 ; ref-Liu-PRB-2007 ; ref-Zeinalipour-Yazdi-JAP-2009 ; ref-Marianetti-PRL-2010 ; ref-Wagner-PRB-2011 ; ref-Mirnezhad-JTS-2012 ; ref-Shao-JCP-2012 , or other theoretical approaches including molecular mechanics ref-Lu-IJAM-2009 ; ref-Fthenakis-JoP-2017 ; ref-Genoese-FrontMat-2019 , and a combination of continuum elasticity theory with other methods ref-Cadelano-PRL-2009 ; ref-Samadikhah-CompMatSci-2012 ; ref-Wei-NanoLet-2013 .
@@
The temperature dependence of various elastic properties of graphene has been examined using MD ref-Shen-MatDes-2010 ; ref-Zhao-JAP-2010 ; ref-deSousa-Intech-2021 , Monte Carlo atomistic simulations ref-Zakharchenko-PRL-2009 , density functional theory ref-Shao-JCP-2012 , or the asymptotic homogenization method ref-Tahani-AMM-2024 . Moreover, a few MD studies have investigated the variation of bond lengths and bond angles with uniaxial tensile loading ref-Zhao-NanoLet-2009 ; ref-Kalosakas-Mat-2022 . The latter one shows results from first principles methods too, and presents analytical expressions for the dependence of these structural parameters on strain ref-Kalosakas-Mat-2022 . The variation of bond lengths and angles with biaxial strain has been given in ref-Dheeraj-Nanotech-2020 . To our knowledge, the effects of temperature and strain on the distributions of bond lengths and bond angles have not yet been presented.
@@
In the present work, we use MD simulations to study the behavior of planar graphene under uniaxial tensile load, considering the influence of temperature. In particular, we implement symplectic integration methods for simulating the system’s time evolution, which allow highly accurate computations for arbitrarily long times. We calculate stress-strain curves at various temperatures and from these results we estimate the variation of several elastic parameters with temperature. Furthermore, we compute bond lengths and bond angles of bulk graphene over a large time-window after thermal equilibrium has been reached, and subsequently analyze these results in order to get the dependence of the corresponding distributions on both stress and temperature. Finally, we present analytical expressions which closely match the numerically obtained distributions of bond lengths and angles. Thus, we describe the detailed dependence of bond lengths and bond angles in graphene on both the applied tensile stress and temperature.
@@ 2 Model and numerical methods
We consider a two-dimensional (2D) model of graphene as a hexagonal lattice of carbon atoms within a plane. Figure 1 illustrates a part of this structure at equilibrium, where the distance between any two neighboring atoms is $r_{0}=1.42$ and the angle formed by three consecutive atoms is $\phi_{0}=2\pi/3$ rad. Furthermore, all carbon atoms have mass $m=12$ amu. In the orientation depicted in Figure 1, the top and bottom edges represent the "armchair edges", while the left and right edges correspond to the "zigzag edges". Furthermore, it is common to call "armchair direction" the horizontal direction in Figure 1, and "zigzag direction" the vertical one.
Figure 1: A schematic of the hexagonal graphene lattice containing $N=42$ atoms, arranged in $N_{I}=6$ columns and $N_{J}=7$ rows. Atoms in column $i$ and row $j$ are indicated in blue and orange, respectively, and the $(i,j)$th atom is colored in gray. The $A$ and $Z$ type bonds, and similarly the $\alpha$ and $\zeta$ type angles, are respectively indicated in red and green (see Section 3.3 for more details on these distinctions).
We refer to columns and rows within the graphene sheet, as indicated by the blue and orange colored atoms, respectively, in Figure 1. A lattice of size $N=N_{I}\times N_{J}$ consists of $N_{I}$ columns, which are indexed by $i$, and $N_{J}$ rows, which are indexed by $j$. The $(i,j)$th atom is indicated in gray in Figure 1. Thus, the configuration shown in this figure corresponds to a $N_{I}\times N_{J}=6\times 7$ lattice, totaling $N=42$ atoms. In our simulations we consider a lattice of $N_{I}=86$ columns and $N_{J}=87$ rows of atoms, resulting in a $N=86\times 87=7482$ total number of carbon atoms. This lattice size is sufficiently large to negate small-size effects ref-Kalosakas-JAP-2013 , allowing thus to represent results on bulk graphene, while being small enough to permit extensive numerical simulations within reasonable CPU times.
A Hamiltonian formalism is used to investigate the in-plane dynamics of the lattice in a similar fashion as that of Ref. ref-Hillebrand-Chaos-2020 . The atomistic force fields describing bond stretching and angle bending deformations are determined through fittings with relevant density functional energy computations ref-Kalosakas-JAP-2013 . In particular, the potential energy of a covalent bond between neighboring atoms at distance $r$ is given by the Morse expression,
@@
where $D=5.7$ eV is the depth of the potential and $a=1.96$-1 is the inverse characteristic length scale of the potential. The angle bending energy term describing a bond angle $\phi$ formed by three consecutive atoms is
@@
where $d=7.0$ eV/rad2 and $d^{\prime}=4$ eV/rad3 are the quadratic and nonlinear coefficients of the potential, respectively.
@@
The total energy of the system (i.e., the values of the model’s Hamiltonian $H$) is the sum of the above potential energy terms for all bond lengths of neighboring atoms and all bond angles between adjacent bonds, and the kinetic energy of each atom. Denoting the total potential energy at time $t$ by $E_{V}(t)$ and the total kinetic energy at $t$ by $E_{K}(t)$, the Hamiltonian
$H(t)=E_{K}(t)+E_{V}(t),$ (3)
@@
is expressed through the $(x(t),y(t))$ positions and the corresponding conjugate momenta of all carbon atoms of the sheet. The time evolution of each atom’s position is governed by the system’s Hamilton’s equations of motion, which conserve the total energy (3).
@@
To study the effects of uniaxial tensile load, a constant force $f$ is applied to all atoms on the appropriate edges ref-Kalosakas-JAP-2013 ; ref-Kalosakas-Mat-2022 : For stress/strain along the armchair direction, $\sigma_{a}$/$\epsilon^{a}$, the force $f$ is applied on the atoms of the zigzag edges, where on the opposite edges opposite forces, directed outwards, are applied. Similarly, for stress/strain along the zigzag direction, $\sigma_{z}$/$\epsilon^{z}$, the force $f$ is applied to the armchair edges, again with opposite forces on opposite edges. Tensile loading leads to additional terms in the Hamiltonian, concerning the relevant edge atoms where the force $f$ is applied. For constant forces, as in our case here, the conservation of the system’s total energy still holds.
@@
In two-dimensional materials like graphene, the stress is given by the force per unit length. Taking into account the distance between successive atoms at the relevant edges where the force is applied, i.e. the atoms of an edge column (row) in Figure 1 concerning stress along the armchair (zigzag) direction, the following relations connect nominal stresses and forces
$\sigma_{a}=\frac{f}{r_{0}\sin(\phi_{0}/2)}\,\,\,\,\mbox{and}\,\,\,\,\sigma_{z}=\frac{f}{0.5\,r_{0}\left[1+\cos(\phi_{0}/2)\right]}.$ (4)
where $r_{0}$ and $\phi_{0}$ are the equilibrium values mentioned above.
@@
To determine at zero temperature the relaxed state of the strained graphene for various applied stresses in any direction, a friction term proportional to the velocity of each atom is incorporated in the MD simulations, setting the friction proportionality coefficient to $\gamma=0.1$ ps-1 (see Ref. ref-Kalosakas-JAP-2013 ). Then, the fourth order Runge-Kutta numerical integration technique is used with an integration time step of $dt=0.005~{}t_{u}$, where $t_{u}=0.0102$ ps represents the time unit for our model. This time step ensures that the relative energy error $Err(t)=\left|H(t)-H(0)\right|/\left|H(0)\right|$ is below $10^{-7}$ in corresponding energy conserving simulations (without the friction term). However, we now simulate the dynamics of the dissipative version of the system until times $t_{f}=3\times 10^{3}~{}t_{u}$, when the total kinetic energy is practically zero ($E_{K}(t_{f})<10^{-16}$ eV). In this way we determine the relaxed equilibrium positions of the atoms for each considered stress $\sigma$.
@@
Based on this equilibrium data of graphene subjected to tensile loads, we embark on the main phase of the numerical investigation: modeling the dynamics of the lattice for different values of stress $\sigma$, at various temperatures $T$, for a long enough time to allow deductions about the thermal equilibrium properties of the stressed material. For these numerical simulations we implement the symplectic integrator ABA864 ref-Blanes-ANM-2013 , with an integration time step $dt=0.06~{}t_{u}$, which results in $Err(t)<10^{-7}$ for all times. This particular integration scheme has been shown to perform very well in balancing computational speed and numerical accuracy for multidimensional Hamiltonian lattices ref-Danieli-MINE-2019 , and was successfully used for examining chaos in graphene ref-Hillebrand-Chaos-2020 .
@@
The relaxed equilibrium positions that have previously been determined for the given value of stress $\sigma$ correspond to a graphene sheet being at zero temperature, without thermal fluctuations. In order to simulate the system at finite temperatures, following a suitable energy-temperature calibration (see subsection 3.1 below), we insert an additional energy density (average energy per site) $e_{N}$ on the relaxed $T=0$ K state. This additional energy is initially provided as solely potential energy, in the form of small random displacements of each atom from the relaxed zero temperature positions. Then these displacements are properly scaled in order to adjust the added energy density $e_{N}$ to the desired value. During evolution the initial potential energy gets shared into kinetic and potential energy and finally the system equilibrates.
@@
In general, for the numerical results presented below we consider 10-20 different individual realizations of the randomly added initial energy, but we have selectively checked the robustness of the data when more realizations are used. For each realization we calculate the temporal evolution of the various quantities of interest, and then compute the average of these time-series over the different realizations in order to obtain the time dependence of the considered quantities for the ensemble. We denote the averaged quantity over the different realizations with angled brackets, e.g. $\left\langle M(t)\right\rangle$ for the measurement of the quantity $M(t)$. We may further determine the average of a thermally equilibrated quantity over time. In such a case we average both over initial realizations and over time intervals, and we denote the computed average by using both an overline and angled brackets, e.g. $\overline{\left\langle M\right\rangle}$ for a variable $M$ at thermal equilibrium.
@@
At finite temperatures the size of graphene sheets exhibits oscillations around their $T=0$ K relaxed configurations due to the thermal energy of the system (discussed further in Section 3.2 below). In order to collect data over sufficiently many such sheet oscillations, we follow the system’s time evolution up to $4\times 10^{3}~{}t_{u}$. Thus, the recording window for all subsequent measurements is from $1\times 10^{3}$ to $4\times 10^{3}~{}t_{u}$, totaling $3000~{}t_{u}$. We have checked the insensitivity of the obtained results on the length of the recording window by testing longer time windows.

### v4

@@ Abstract
Using molecular dynamics simulations in a planar graphene sheet, we investigate the temperature dependence of its mechanical behavior under uniaxial tensile stress applied either along the armchair or the zigzag direction. Stress-strain curves are calculated for different temperatures and the corresponding dependence of various elastic parameters, is discussed. Fracture stress and strain, as well as the Young modulus, decrease almost linearly with temperature, in accordance with previous investigations. An almost linear variation of the third-order elastic modulus with temperature is demonstrated, revealing opposite trends for uniaxial loadings along the armchair or the zigzag direction. The detailed dependence of the distributions of bond lengths and bond angles both on strain and temperature is presented for the first time, along with approximate analytical expressions. The latter describe accurately the numerically obtained distributions.
@@ 1 Introduction
Since the discovery of graphene, there have been a number of investigations into its mechanical behavior. Despite the difficulty of applying controlled mechanical loads at the nanoscale, experimental studies have verified an exceptional value of stiffness and extremely high tensile strength ref-Lee-Science-2008 ; ref-Cao-NatComm-2020 ; ref-Varillas-MechSci-2024 , in accordance with corresponding theoretical predictions. For example, in Reference ref-Lee-Science-2008 the mechanical properties of graphene were deduced from nanoindentation experiments on circularly clamped samples, while in References ref-Cao-NatComm-2020 ; ref-Varillas-MechSci-2024 push-to-pull testing was employed to directly apply uniaxial strain. There exist a number of related numerical investigations using molecular dynamics (MD) simulations with a variety of potential functions ref-Zakharchenko-PRL-2009 ; ref-Zhao-NanoLet-2009 ; ref-Neek-Amal-PRB-2010 ; ref-Shen-MatDes-2010 ; ref-Tsai-MatDes-2010 ; ref-Zhao-JAP-2010 ; ref-Wei-JCP-2011 ; ref-Dmitriev-CompMatSci-2012 ; ref-Kalosakas-JAP-2013 ; ref-Baimova-EUPhyLet-2013 ; ref-Zhang-CompMatSci-2013 ; ref-Sgouros-2DMat-2016 ; ref-Sgouros-SciRep-2018 ; ref-Lebedeva-PhysE-2019 ; ref-Savin-MechofMat-2019 ; ref-Aparicio-Nanotech-2021 ; ref-deSousa-Intech-2021 ; ref-Kalosakas-Mat-2021 ; ref-Flores-PhysChem-2022 ; ref-Fthenakis-FrontChem-2022 ; ref-Varma-IntJMechMatDes-2022 ; ref-Pacheco-Sanjuan-Carbon-2023 ; ref-Sgouros-PRE-2024 , density functional theory ref-Kudin-PRB-2001 ; ref-Liu-PRB-2007 ; ref-Zeinalipour-Yazdi-JAP-2009 ; ref-Marianetti-PRL-2010 ; ref-Wagner-PRB-2011 ; ref-Mirnezhad-JTS-2012 ; ref-Shao-JCP-2012 , or other theoretical approaches including molecular mechanics ref-Lu-IJAM-2009 ; ref-Fthenakis-JoP-2017 ; ref-Genoese-FrontMat-2019 , and combinations of continuum elasticity theory with other methods ref-Cadelano-PRL-2009 ; ref-Samadikhah-CompMatSci-2012 ; ref-Wei-NanoLet-2013 ; ref-Tanhadoust-DiamRelMat-2022 .
@@
The temperature dependence of various elastic properties of graphene has been examined using MD ref-Shen-MatDes-2010 ; ref-Zhao-JAP-2010 ; ref-deSousa-Intech-2021 , Monte Carlo atomistic simulations ref-Zakharchenko-PRL-2009 , density functional theory ref-Shao-JCP-2012 , or the asymptotic homogenization method ref-Tahani-AMM-2024 . Moreover, a few MD studies have investigated the variation of bond lengths and bond angles with uniaxial tensile loading ref-Zhao-NanoLet-2009 ; ref-Kalosakas-Mat-2022 . The latter work shows results from first principles methods too, and presents analytical expressions for the dependence of these structural parameters on strain ref-Kalosakas-Mat-2022 . The variation of bond lengths and angles with biaxial strain has been discussed in ref-Dheeraj-Nanotech-2020 . To our knowledge, the dependence of the distributions of bond lengths and bond angles on both temperature and strain has not yet been examined.
@@
In the present work, we use MD simulations to study the behavior of planar graphene under uniaxial tensile load, considering the influence of temperature. In particular, we implement symplectic integration methods for simulating the system’s time evolution, which allow highly accurate computations for arbitrarily long times. We calculate stress-strain curves at various temperatures and from these results we estimate the variation of several elastic parameters with temperature. Our findings for the Young modulus, fracture strength, and failure strain, are in agreement with previous studies. The temperature dependence of the the third-order elastic modulus has not been presented up to now.
Furthermore, we compute bond lengths and bond angles of bulk graphene over a large time-window after thermal equilibrium has been reached, and subsequently analyze these results in order to get the dependence of the corresponding distributions on both stress and temperature. Finally, we present analytical expressions which closely match the numerically obtained distributions of bond lengths and angles. Thus, we describe, for the first time to the best of our knowledge, the detailed dependence of these structural properties of graphene on both the applied tensile stress and temperature.
@@ 2 Model and numerical methods
We consider a two-dimensional (2D) model of graphene as a hexagonal lattice of carbon atoms within a plane. Figure 1 illustrates a part of this structure at equilibrium, where the distance between any two neighboring atoms is $r_{0}=1.42$ Å and the angle formed by three consecutive atoms is $\phi_{0}=2\pi/3$ rad. Furthermore, all carbon atoms have mass $m=12$ amu. In the orientation depicted in Figure 1, the top and bottom edges represent the "armchair edges", while the left and right edges correspond to the "zigzag edges". Furthermore, it is common to call "armchair direction" the horizontal direction in Figure 1, and "zigzag direction" the vertical one.
Figure 1: A schematic of the hexagonal graphene lattice depicting $N=42$ atoms, arranged in $N_{I}=6$ columns and $N_{J}=7$ rows. Atoms in column $i$ and row $j$ are indicated in blue and orange, respectively, and the $(i,j)$th atom is colored in gray. The AA and $Z$ type bonds, and similarly the $\alpha$ and $\zeta$ type angles, are respectively indicated in red and green (see Section 3.3 for more details on these distinctions).
We refer to columns and rows within the graphene sheet, as indicated by the blue and orange colored atoms, respectively, in Figure 1. A lattice of size $N=N_{I}\times N_{J}$ consists of $N_{I}$ columns, which are indexed by $i$, and $N_{J}$ rows, which are indexed by $j$. The $(i,j)$th atom is indicated in gray in Figure 1. Thus, the configuration shown in this figure corresponds to a $N_{I}\times N_{J}=6\times 7$ lattice, totaling $N=42$ atoms. In our simulations we consider a lattice of $N_{I}=86$ columns and $N_{J}=87$ rows of atoms, resulting in a $N=86\times 87=7482$ total number of carbon atoms. This lattice size is sufficiently large to negate small-size effects ref-Kalosakas-JAP-2013 , allowing thus to represent the behavior of bulk graphene, while still permitting extensive numerical simulations within reasonable CPU times. We have further confirmed that indicative data for larger lattices are similar to the ones presented here.
A Hamiltonian formalism is used to investigate the in-plane dynamics of the lattice in a similar fashion as that of Reference ref-Hillebrand-Chaos-2020 . The atomistic force field, describing bond stretching and angle bending deformations, has been determined through fittings with relevant density functional energy computations ref-Kalosakas-JAP-2013 . In particular, the potential energy of a covalent bond between nearest neighboring carbon atoms at distance $r$ is given by the Morse expression,
@@
where $D=5.7$ eV is the depth of the potential well and $a=1.96$ Å-1 is the inverse characteristic length scale of the potential. The angle bending energy term describing a bond angle $\phi$ formed by three consecutive atoms is
@@
where $d=7.0$ eV/rad2 and $d^{\prime}=4$ eV/rad3 are the constants of the quadratic and nonlinear term of the potential, respectively.
A limitation of the used force field, given by Equations (1) and (2), is the decoupling of the bond stretching and angle bending variables. Though this is a standard approximation in atomistic simulations, in real systems these degrees of freedom are coupled. However, the predictions of this model for the Young modulus and the intrinsic strength of graphene are in good agreement with the experimental estimates of Reference ref-Lee-Science-2008 , while the stress-strain response is in accordance with corresponding calculations from first principles ref-Kalosakas-JAP-2013 . A more drastic approximation is that the out-of-plane atomic displacements are ignored. At finite temperatures ripples and other out-of-plane deformations spontaneously appear in graphene due to thermal fluctuations ref-Fasolino-NMat-2007 ; ref-Meyer-Nature-2007 . However, we expect that at relatively low temperatures these non-planar deformations would be suppressed when graphene is uniaxially stressed. Stretching has been experimentally used to flatten graphene ref-Cao-NatComm-2020 and this has been further confirmed by MD simulations ref-Lee-NanoResLets-2015 . Thus, we have considered here only temperatures up to 700 K and not higher ones, even though graphene exists at much larger temperatures. As will be discussed later, the intrinsic strength and fracture strain obtained via our model at room temperature are in good agreement with those obtained from fully three-dimensional MD simulations employing different interatomic potentials ref-Zhao-NanoLet-2009 ; ref-Lee-NanoResLets-2015 ; ref-Varillas-MechSci-2024 .
@@
The total energy of the system (i.e., the values of the model’s Hamiltonian $H$) is the sum of the above potential energy terms for all bond lengths between nearest neighboring atoms and all bond angles between adjacent bonds, and the kinetic energy of each atom. Denoting the total potential energy at time $t$ by $E_{V}(t)$ and the total kinetic energy at $t$ by $E_{K}(t)$, the Hamiltonian
$H=E_{K}(t)+E_{V}(t),$ (3)
@@
is expressed through the positions $(x(t),y(t))$ and the corresponding conjugate momenta of all carbon atoms within the considered graphene sheet. The time evolution of each atom’s position and momentum is governed by the system’s Hamilton’s equations of motion, which conserve the total energy, Equation (3).
@@
To investigate the effects of uniaxial tensile load, a constant force $f$ is applied to all atoms on the appropriate edges of the sheet ref-Kalosakas-JAP-2013 ; ref-Kalosakas-Mat-2022 : For stress/strain along the armchair direction, the force $f$ is applied on the atoms of the zigzag edges, where on the opposite edges opposite forces, directed outwards, are applied. Similarly, for stress/strain along the zigzag direction, the force $f$ is applied to the armchair edges, again with opposite forces on opposite edges. Tensile loading results in additional terms in the Hamiltonian, given by appropriate products of the relevant edge atom displacements with the applied force $f$. For constant forces, as in our case here, the conservation of the system’s total energy still holds.
We emphasize that we perform stress-controlled simulations here, where we fix the forces (stress) and directly compute the resulting strain. This is a natural choice in MD, in contrast to imposing fixed displacements (strain-controlled simulations) which is preferred in first principles’ studies. In principle, these two methods of strain- or stress-controlled simulations are equivalent for estimating the stress-strain mechanical response of the system. For example, one can see in Figures 4 and 5 of Reference ref-Kalosakas-JAP-2013 the direct comparison of stress-controlled MD data and strain-controlled density functional theory data, which produce identical results at least in the linear response regime.
@@
In two-dimensional materials like graphene, the stress is expressed as force per unit length. Taking into account the distance between successive atoms at the relevant edges where the force is applied, i.e. the nearest neighboring atoms along an edge column, or row, in Figure 1 concerning stress in the armchair, or zigzag respectively, direction, the following relations connect nominal stresses and forces
$\sigma_{a}=\frac{f}{r_{0}\sin(\phi_{0}/2)}\,\,\,\,\mbox{and}\,\,\,\,\sigma_{z}=\frac{f}{0.5\,r_{0}\left[1+\cos(\phi_{0}/2)\right]},$ (4)
where the indices $a$ and $z$ denote stress in the armchair and zigzag direction, respectively, while $r_{0}$ and $\phi_{0}$ are the equilibrium values mentioned above.
@@
To determine at zero temperature the relaxed state of the strained graphene for various applied stresses in any direction, a friction term proportional to the velocity of each atom is incorporated in the MD simulations, setting the friction coefficient to $\gamma=0.1$ ps-1; see Reference ref-Kalosakas-JAP-2013 and the discussion therein. Then, the fourth order Runge-Kutta numerical integration technique is used with an integration time step of $dt=0.005~t_{u}$, where $t_{u}=0.0102$ ps represents the time unit in our model. This time step ensures that the relative energy error $Err(t)=\left|H(t)-H(0)\right|/\left|H(0)\right|$ is below $10^{-7}$ in corresponding energy conserving simulations where the friction term is absent. However, we now simulate the dynamics of the dissipative version of the system until times $t_{f}=3\times 10^{3}~t_{u}$, when the total kinetic energy is practically zero ($E_{K}(t_{f})<10^{-16}$ eV). In this way we determine the relaxed equilibrium positions of the atoms within the lattice, for each considered value of stress $\sigma$.
@@
Based on this equilibrium configuration of graphene subjected to tensile loads, we embark on the main phase of the numerical investigation: following the dynamics of the lattice for a fixed value of stress $\sigma$, at various temperatures $T$, for a long enough time to allow deductions about the thermal equilibrium properties of the stressed material. For these numerical simulations we implement the symplectic integrator ABA864 ref-Blanes-ANM-2013 , with an integration time step $dt=0.06~t_{u}$, which results in relative errors of the total energy $Err(t)<10^{-7}$ for all times. This particular integration scheme has been shown to perform very well in balancing computational speed and numerical accuracy for multidimensional Hamiltonian lattices ref-Danieli-MINE-2019 , and was successfully used for examining chaos in graphene ref-Hillebrand-Chaos-2020 .
@@
The relaxed equilibrium positions that have previously been determined for the given value of stress $\sigma$ correspond to a graphene sheet being at zero temperature, without thermal fluctuations. In order to simulate the system at finite temperatures, following a suitable energy-temperature calibration (see subsection 3.1 below), we randomly insert an additional energy density (average energy per site) $e_{N}$ on the relaxed $T=0$ K state. This additional energy is initially provided as solely potential energy, in the form of small random displacements of each atom from the relaxed zero temperature positions. Then these displacements are properly scaled in order to adjust the added energy density $e_{N}$ to the desired value corresponding to the simulated temperature. During evolution the initial potential energy gets shared into kinetic and potential energy and eventually the system equilibrates.
@@
In general, for the numerical results presented in the next section we consider 10-20 different individual realizations of the randomly added initial energy, but we have selectively checked the robustness of the data when more realizations are used. For each realization we calculate the temporal evolution of the various quantities of interest, and then compute the average of these time-series over the different realizations in order to obtain the time dependence of the considered quantities for the ensemble. We denote such an averaged quantity over the different realizations with angled brackets, e.g. $\left\langle M(t)\right\rangle$ for the measurement of the quantity $M(t)$. We may further determine the average of a thermally equilibrated quantity over time. In such a case we average both over initial realizations and over time intervals, and we denote the computed average by using both an overline and angled brackets, e.g. $\overline{\left\langle M\right\rangle}$ for a variable $M$ at thermal equilibrium.
@@
At finite temperatures the size of graphene sheets exhibits oscillations around their $T=0$ K relaxed configurations due to the thermal energy of the system (discussed further in Section 3.2 below). In order to collect data over sufficiently many such sheet oscillations, we follow the system’s time evolution up to $4\times 10^{3}~t_{u}$. The recording window for all subsequent measurements is from $1\times 10^{3}$ to $4\times 10^{3}~t_{u}$, totaling $3000~t_{u}$. We have checked the insensitivity of the obtained results on the length of the recording window by testing longer time windows.